The curve x^4+y^4=4a^2xy is symmetric with respect to the origin.
True or false with full explanation
1
Expert's answer
2020-11-23T19:32:38-0500

Let "(x_0, y_0)" is the point of the equation, that is "x_0^4+y_0^4=4a^2x_0y_0."

Consider "(-x_0, -y_0)" – symmetric point with respect to the origin.

Substitute it into the equation:

"(-x_0)^4+(-y_0)^4=4a^2(-x_0)(-y_0),"

or "x_0^4+y_0^4=4a^2x_0y_0."

So the curve is symmetric with respect to the origin.